The perimeter of a parallelogram is 36 cm.if a diagonal is 6cm less its length but 2 cm more than its width,what are the dimensions of its parallelogram?
Answer by Gogonati(855)

Since the opposite sides of the parallelogram are equal and the perimeter is 36, half of the perimeter is 18 cm, and express the sum of two consecutive sides of the parallelogram.
Lets the width = x cm then the length will be 18-x cm. Since the diagonal is 6 cm less its length we write: d=18-x-6=12-x cm. Since the diagonal is 2 more than its width we write: d=x+2 cm.Now we can write:x+2=12-x, solving this equation we get x=5 cm.
Answer:The dimensions of the parallelogram are: width 5 cm and the length 18-5=13 cm.